# Convexified Message-Passing Graph Neural Networks

### v1

@@ B.3 Valid Activation Functions
Appendix C Rephrasing the Nonlinear Filter at each Layer
@@
=${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle.$ (18)
Defining the matrix $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k+1}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$, ∎
Appendix D Kernelization for Nonlinear Activations
In this section, we elaborate on the kernelization for the case of nonlinear activation functions, briefly discussed in Section 4.2. Recall that $\sigma$ denotes, for brevity, its entrywise application in (2). Further, the of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$’s output is given by $\sigma(\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle)$ by (18). According to Appendix B, consider a sufficiently smooth valid activation function $\sigma$, and a proper choice of kernel $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $[\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem [44, Theorem 4.49] implies that there exists a feature mapping $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$:
@@ Appendix D Kernelization for Nonlinear Activations
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. By Corollary B.1 and Corollary B.3, we have that $\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}\leq C_{\sigma,\ell}(\|{\mathbf{a}}^{f}_{\ell k}\|_{2})$, provided a monotonically increasing $C_{\sigma}$ which depends on the kernel $\kappa_{\ell}$. Accordingly, $\varphi_{\ell}(\mathbf{z})$ may be utilized as the vectorized representation of each output features $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$, where $\bar{\mathbf{a}}^{f}_{\ell k}$ constitutes the linear graph filters, thus reducing the problem to training a GCN with the identity activation function.
@@
Accordingly, replacing the vector $\mathbf{\beta}_{\ell k}^{f}$ on the right-hand side of (23) by $((Q_{\ell}^{k})^{\top})^{\dagger}(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ won’t affect the empirical risk. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for $k\in[K]\cup\{0\}$, we thereby propose the following parameterization function:
@@
Consequently, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ is the matrix whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under the reparameterization in (25) which satisfy the following constraints:
@@
Appendix E More Details on Matrix Approximation and Factorization Methods
As mentioned in Section 4.4, the time complexity of Algorithm 2 largely depends on the width $P$ of each factorization matrix $Q_{\ell}^{k}$ and the method used to obtain it (line 3). A naive choice of $P=nm$ corresponds to solving the full kernelized problem, which can be computationally expensive. For instance, using Cholesky factorization with $P=nm$ [8] incurs a complexity of $\mathcal{O}(m^{3}n^{3})$ and space complexity of $\mathcal{O}(m^{2}n^{2})$.
However, to gain scalability, we use Nyström approximation [48]: an approximation $\mathcal{K}_{\ell}^{k}\approx Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$ for $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ is obtained by randomly sampling $P$ rows/columns from the original $\mathcal{K}_{\ell}^{k}$, which takes $\mathcal{O}(P^{2}nm)$ time. In practice, it randomly samples training examples, computes their kernel matrix and represents each training example by its similarity to the sampled examples and kernel matrix. It is well-known that this method can significantly accelerate our computations [48, 9], which is attained by using the approximate kernel matrix instead of the full one. Another advantage is that it is not necessary to compute or store the full kernel matrix, but only a submatrix requiring space complexity of $\mathcal{O}(Pmn)$.
@@ Appendix E More Details on Matrix Approximation and Factorization Methods
Another possible approximation method is the random feature approximation [35], which can be executed in $\mathcal{O}(nmP\log F_{\ell})$ time [26]. Overall, using approximations significantly accelerates our algorithm’s performance in terms of time and space complexity.
Appendix F Proof of Theorem 4.7
@@
Consider models with a single hidden layer for binary classification (i.e., $L=G=1$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices B.1-B.2). For any valid activation function $\sigma$ (Recall Appendix B.3), there is a constant $C_{\sigma,L}(\mathcal{B}_{L})$ such that with the radius $R_{L}:=C_{\sigma,L}(\mathcal{B}_{L})\sqrt{F}$, the expected generalization error is at most:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]+\frac{cMR_{L}\sqrt{\log(n(K+1))\sum_{k=0}^{K}\mathbb{E}[||K_{L}^{k}(\mathcal{X})||_{2}]}}{\sqrt{m}}$ (27)
@@
We begin by proving several properties that apply to the general multi-class case with models comprising of multiple hidden layers, and then we specialize them to a single-hidden-layer models in the binary classification case. First, a wealthier function class shall be considered, which contains the class of GCNs. For each layer $\ell$, let $\mathcal{H}_{\ell}$ be the RKHS induced by the kernel function $\kappa_{\ell}$ and $||\cdot||_{\mathcal{H}_{\ell}}$ be the associated Hilbert norm. Recall that, given any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$, we denote its $(k+1)^{\text{th}}$ row as $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$. Using this notation, consider the following function class:
$\mathcal{F}_{\text{cgcn}}:=\{\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}:\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}\quad\forall\ell\in[L]\}\text{, where}$ (28)
$\mathcal{F}_{\ell,cgcn}:=\Bigg{\{}\phi^{{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}^{\star}}\rightarrow\mathbb{R}^{n\times F_{\ell}}:F_{\ell-1}^{\star}<\infty\text{ and }[\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}))$
[…]
@@
Recall that that any GCN $\Phi\in\mathcal{F}_{\text{gcn}}$ induces a nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ which can be characterized by $\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle)$, i.e., the GCN $\Phi$ is given by the following composition of functions $\Phi=\tau_{L}\circ\dots\circ\tau_{1}$, where we defined $\tau_{\ell}:\mathbf{z}\mapsto(\tau_{\ell}^{f}(\mathbf{z}))_{f\in[F_{\ell}]}$. Given any valid activation function, Corollary B.1 and Corollary B.3 yield that $\tau_{\ell}^{f}$ resides within the reproducing kernel Hilbert space $\mathcal{H}_{\ell}$ and its Hilbert norm is upper bounded as $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\leq C_{\sigma}(\|\sum_{k=0}^{K}\mathbf{a}^{f}_{\ell k}\|_{2})\leq C_{\sigma}(R_{\ell})\sqrt{K+1}$, which clearly satisfies the constraint in (28) for any layer $\ell$, as desired. ∎
Recalling that $\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\cdot)$ is the predictor at each layer $\ell$ produced by Algorithm 2, we now formally prove that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer is an empirical risk minimizer in the function class $\mathcal{F}_{\text{cgcn}}$.
Lemma F.2.
@@
:=$\mathcal{F}_{\mathcal{B}_{\ell}}:=\Bigg{\{}\Phi^{\mathcal{A}}=\Psi^{\mathcal{A}_{L}}\circ\dots\circ\Psi^{\mathcal{A}_{1}}\Bigg{|}$ $\Psi^{\mathcal{A}_{\ell}}:\mathcal{X}\mapsto\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}\text{ and }$ (31)
$\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}\text{ and }\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\Bigg{\}}$
@@
=${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\Bigg{\langle}\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\Bigg{\rangle}$ (33)
Let $\mathbf{v}_{\ell}^{k}(\mathbf{z})\in\mathbb{R}^{nm}$ be a vector whose $(i,j)$-th coordinate equals to $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$. Similarly to Appendix D, denote $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$, note that $\tau_{\ell}^{f}$ can be rephrased as $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle\mathbf{v}_{\ell}^{k}(\mathbf{z}),((Q_{\ell}^{k})^{\top})^{\dagger}\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. Thus, (22) implies that the Hilbert norm of the function $\tau_{\ell}^{f}$ satisfies:

### v2

@@ C.3 Valid Activation Functions
Appendix D REPHRASING THE NONLINEAR FILTER AT EACH LAYER
@@
${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle.$ (18)
Appendix E KERNELIZATION FOR NONLINEAR ACTIVATIONS
In this section, we elaborate on the kernelization for the case of nonlinear activation functions, briefly discussed in Section 4.2. Recall that $\sigma$ denotes, for brevity, its entrywise application in (2). Further, the of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$’s output is given by $\sigma(\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle)$ by (18). According to Appendix C, consider a sufficiently smooth valid activation function $\sigma$, and a proper choice of kernel $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem (Steinwart, 2008, Theorem 4.49) implies that there exists a feature mapping $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$:
@@ Appendix E KERNELIZATION FOR NONLINEAR ACTIVATIONS
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. By Corollary C.1 and Corollary C.3, we have that $\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}\leq C_{\sigma,\ell}(\|{\mathbf{a}}^{f}_{\ell k}\|_{2})$, provided a monotonically increasing $C_{\sigma}$ which depends on the kernel $\kappa_{\ell}$. Accordingly, $\varphi_{\ell}(\mathbf{z})$ may be utilized as the vectorized representation of each output features $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$, where $\bar{\mathbf{a}}^{f}_{\ell k}$ constitutes the linear graph filters, thus reducing the problem to training a GCN with the identity activation function.
@@
Accordingly, replacing the vector $\mathbf{\beta}_{\ell k}^{f}$ on the right-hand side of (23) by $((Q_{\ell}^{k})^{\top})^{\dagger}(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ won’t affect the empirical risk. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for $k\in[K]\cup\{0\}$, we thereby propose the following parameterization function:
@@
Consequently, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ is the matrix whose $(k+1)^{\text{th}}$ row is $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under the reparameterization in (25) which satisfy the following constraints:
@@
Appendix F MORE DETAILS ON MATRIX APPROXIMATION AND FACTORIZATION METHODS
As mentioned in Section 4.4, the time complexity of Algorithm 2 largely depends on the width $P$ of each factorization matrix $Q_{\ell}^{k}$ and the method used to obtain it (line 3). A naive choice of $P=nm$ corresponds to solving the full kernelized problem, which can be computationally expensive. For instance, using Cholesky factorization with $P=nm$ (Dereniowski and Kubale, 2003) incurs a complexity of $\mathcal{O}(m^{3}n^{3})$ and space complexity of $\mathcal{O}(m^{2}n^{2})$.
However, to gain scalability, we use Nyström approximation (Williams and Seeger, 2001): an approximation $\mathcal{K}_{\ell}^{k}\approx Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$ for $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ is obtained by randomly sampling $P$ rows/columns from the original $\mathcal{K}_{\ell}^{k}$, which takes $\mathcal{O}(P^{2}nm)$ time. In practice, it randomly samples training examples, computes their kernel matrix and represents each training example by its similarity to the sampled examples and kernel matrix. It is well-known that this method can significantly accelerate our computations (Williams and Seeger, 2001; Drineas and Mahoney, 2005), which is attained by using the approximate kernel matrix instead of the full one. Another advantage is that it is not necessary to compute or store the full kernel matrix, but only a submatrix requiring space complexity of $\mathcal{O}(Pmn)$.
@@ Appendix F MORE DETAILS ON MATRIX APPROXIMATION AND FACTORIZATION METHODS
Another possible approximation method is the random feature approximation (Rahimi and Recht, 2007), which can be executed in $\mathcal{O}(nmP\log F_{\ell})$ time (Le et al., 2013). Overall, using approximations significantly accelerates our algorithm’s performance in terms of time and space complexity.
Appendix G PROOF OF THEOREM 4.7
@@
Consider models with a single hidden layer for binary classification (i.e., $L=1,G=2$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices C.1-C.2). For any valid activation function $\sigma$ (Recall Appendix C.3), there is a constant $C_{\sigma,L}(\mathcal{B}_{L})$ such that with the radius $R_{L}:=C_{\sigma,L}(\mathcal{B}_{L})\sqrt{2F}$, the expected generalization error is at most:
[…]
@@
We begin by proving several properties that apply to the general multi-class case with models comprising of multiple hidden layers, and then we specialize them to single-hidden-layer models in the binary classification case. First, a wealthier function class shall be considered, which contains the class of GCNs. For each layer $\ell$, let $\mathcal{H}_{\ell}$ be the RKHS induced by the kernel function $\kappa_{\ell}$ and $||\cdot||_{\mathcal{H}_{\ell}}$ be the associated Hilbert norm. Recall that, given any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$, we denote its $(k+1)^{\text{th}}$ row as $\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$. Using this notation, consider the following function class:
$\mathcal{F}_{\text{cgcn}}=\{\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}:\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}\quad\forall\ell\in[L]\}\text{, where}$ (28)
$\mathcal{F}_{\ell,cgcn}=\Bigg\{\phi^{{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}^{\star}}\rightarrow\mathbb{R}^{n\times F_{\ell}}:F_{\ell-1}^{\star}<\infty\text{ and }[\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}))$
$\text{ where }\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}^{\star}}[\mathbf{z}_{k}]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}\text{ and }$
@@
Recall that that any GCN $\Phi\in\mathcal{F}_{\text{gcn}}$ induces a nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ which can be characterized by $\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle)$, i.e., the GCN $\Phi$ is given by the following composition of functions $\Phi=\tau_{L}\circ\dots\circ\tau_{1}$, where we defined $\tau_{\ell}:\mathbf{z}\mapsto(\tau_{\ell}^{f}(\mathbf{z}))_{f\in[F_{\ell}]}$. Given any valid activation function, Corollary C.1 and Corollary C.3 yield that $\tau_{\ell}^{f}$ resides within the reproducing kernel Hilbert space $\mathcal{H}_{\ell}$ and its Hilbert norm is upper bounded as $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\leq C_{\sigma}(\|\sum_{k=0}^{K}\mathbf{a}^{f}_{\ell k}\|_{2})\leq C_{\sigma}(R_{\ell})\sqrt{K+1}$, which clearly satisfies the constraint in (28) for any layer $\ell$, as desired. ∎
[…]
Lemma G.2.
@@
$\mathcal{F}_{\mathcal{B}_{\ell}}=\Bigg\{\Phi^{\mathcal{A}}=\Psi^{\mathcal{A}_{L}}\circ\dots\circ\Psi^{\mathcal{A}_{1}}\Bigg|$ $\Psi^{\mathcal{A}_{\ell}}:\mathcal{X}\mapsto\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}\text{ and }$ (31)
[…]
@@
${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\Bigg\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\Bigg\rangle$ (33)
Let $\mathbf{v}_{\ell}^{k}(\mathbf{z})\in\mathbb{R}^{nm}$ be a vector whose $(i,j)$-th coordinate equals to $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$. Similarly to Appendix E, denote $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$, note that $\tau_{\ell}^{f}$ can be rephrased as $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle\mathbf{v}_{\ell}^{k}(\mathbf{z}),((Q_{\ell}^{k})^{\top})^{\dagger}\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. Thus, (22) implies that the Hilbert norm of the function $\tau_{\ell}^{f}$ satisfies: